# Probabilistic Kernel Function for Fast Angle Testing

### v2

@@ 7.1 Comparison with CEOs
As demonstrated in Sec. 6.1, the results of CEOs are directly used to accelerate other similarity search processes (Pham & Liu, 2022; Xu & Pham, 2024). In this context, we focus solely on the improvement of CEOs itself. We show that KS1, equipped with the structures $S_{\text{sym}}(m,1)$ and $S_{\text{pol}}(m,1)$, can slightly outperform CEOs($m$) on the original task of CEOs, that is, $k$-MIPS, where $m$ denotes the number of projection vectors and was set to 2048, following the standard configuration of the original CEOs. Since the only difference among the compared approaches is the configuration of the projection vectors, we use a unified algorithm (see construction of projection structure Alg. 4 and MIPS query processing Alg. 5 in Appendix) with the configuration of projection vectors as an input to compare their recall rates. From the results in Table. 1, we observe that: (1) in most cases, KS1 with the two proposed structures achieves slightly better performance than CEOs, supporting our claim that a smaller reference angle yields a more accurate estimation, and (2) $S_{\text{pol}}$ generally achieves a higher recall rate than $S_{\text{sym}}$, verifying that a configuration closer to the best covering yields better performance.
@@ 7.2 ANNS Performance
We chose ScaNN (Guo et al., 2020), HNSW (Malkov & Yashunin, 2020), and HNSW+PEOs (Lu et al., 2024) as baselines, where ScaNN is a state-of-the-art quantization-based approach that performs better than IVFPQFS, and HNSW+PEOs (Lu et al., 2024) is a state-of-the-art graph-based approach that outperforms FINGER (Chen et al., 2023) and Glass (Zilliz, 2023). Like HNSW+PEOs, KS2 is implemented on HNSW, dubbed HNSW+KS2. DiskANN (Subramanya et al., 2019) is excluded because it is orthogonal to KS2 and focuses on optimization for external storage. ADSampling (Gao & Long, 2023) is excluded because it is designed for a non-SIMD environment. The parameter settings of all compared approaches and Additional experimental results can be found in Appendix D.
@@
(2) Query performance. From the results in Fig. 1, we make the following observations. (i) Except for Word, HNSW+KS2 achieves the best performance among all compared methods. In particular, HNSW+KS2 accelerates HNSW by a factor of 2.5 to 3, and is 1.1 to 1.3 times faster than HNSW+PEOs, demonstrating the superiority of KS2 over PEOs. (ii) Compared with ScaNN, the advantage of HNSW+KS2 is especially evident in the recall region below 85%, highlighting the high efficiency of the routing test. On the other hand, in the high-recall region for Word, ScaNN outperforms HSNW+KS2 due to the connectivity issues of HNSW.
@@ Acknowledgements
This work is partially supported by JSPS Kakenhi JP23K17456, JP23K24850, JP23K25157, JP23K28096, JP25H01117, JP25K00161, and JP25K21272, and JST CREST JPMJCR22M2.
@@ Appendix A Notations
$\|\bm{e}\|$ Length of edge $e$
$\tau$ Threshold determined by the result list of graph
@@
$\bm{u}^{i}_{j}$ $j$-th element of the $i$-th sub-vector of $\bm{u}$, where $1\leq i\leq L$ and $1\leq j\leq m$
@@ Appendix B Proof of Lemmas
If $u$ is a vector randomly drawn from $C$, the CDF of $\left\langle\bm{v},\bm{u}\right\rangle$ is $I_{t}(\frac{d-2}{2},\frac{d-2}{2})$, where $t=\frac{1}{2}+\frac{x-\cos\phi\cos\psi}{2\sin\phi\sin\psi}$.
@@
We consider the following four cases based on the value of $\cos\left\langle\bm{q},\bm{v}\right\rangle$.
Case 1: $\bm{0<\cos\left\langle q,v\right\rangle<1}$.
@@
Claim: Let $O,A,B,C$ be four points $\mathbb{R}^{d}$. $\overrightarrow{OA}$ is perpendicular to $\overrightarrow{AB}$, and $\overrightarrow{OA}$ is perpendicular to $\overrightarrow{AC}$. The angle between $\overrightarrow{OA}$ and $\overrightarrow{OC}$ is $\psi$, and the angle between $\overrightarrow{OA}$ and $\overrightarrow{OB}$ is $\phi$. If the angle between $\overrightarrow{AB}$ and $\overrightarrow{AC}$ is $\alpha$, then $\cos\beta$, where $\beta$ denotes the angle between $\overrightarrow{OB}$ and $\overrightarrow{OC}$, can be expressed as follows.
@@
Case 2: $\bm{\cos\left\langle q,v\right\rangle=0}$.
@@
Case 3: $\bm{-1<\cos\left\langle q,v\right\rangle<0}$.
@@
Case 4: $\bm{\cos\left\langle q,v\right\rangle=1}$ or $\bm{\cos\left\langle q,v\right\rangle=-1}$.
In this case, $\cos\left\langle\bm{q},\bm{u}\right\rangle=\pm\cos\psi$, and the conclusion is trivial.
@@
2   Generate $m$ points randomly and independently on $S^{d^{\prime}-1}$  Scale the norm of all $m$ points in this iteration to $1/\sqrt{L}$ and collect the vectors after scaling
@@ B.3 Estimation of $J(S_{\mathrm{sym}}(m,L))$
Then we only need to focus on $J(S_{\text{ran}}(m,L))$ and prove that it is equal to the RHS of Ineq. (29). Let $\bm{v}=[\bm{v_{1}},\bm{v_{2}},\cdots,\bm{v_{L}}]\in\mathbb{R}^{n}$ be a vector drawn randomly from $\mathbb{S}^{d-1}$, where $d$ is assumed to be divisible by $L$ and $d=Ld^{\prime}$. Let $r_{L}(\bm{v})$ be the regularized vector w.r.t. $\bm{v}$. That is,
@@
Let $T(d,L)=\cos\left\langle\bm{v},r_{L}(\bm{v})\right\rangle$. For every $i$, let $\bm{u_{i}}=\bm{v_{i}}/\|\bm{v_{i}}\|\in\mathbb{S}^{d^{\prime}-1}$. Then we select $m$ vectors from $\mathbb{S}^{d^{\prime}-1}$ randomly and independently. Suppose that, among $m$ generated vectors, $\bm{w}$ is the vector having the smallest angle to $\bm{u_{i}}$, and we use $T^{\prime}(d^{\prime},L)$ to denote $\cos\left\langle\bm{w},\bm{u_{i}}\right\rangle$. Since the choice of $w$ for every $u_{i}$ is independent to the value of $\cos\left\langle\bm{v},r_{L}(\bm{v})\right\rangle$, we have the following result:
$\mathbb{E}[A_{S}(\bm{v})]=\mathbb{E}[T(d,L)]\times\mathbb{E}[T^{\prime}(d^{\prime},L)].$ (32)
@@
$E[T(d,L)]=\sqrt{L}\frac{\Gamma(\frac{d+L}{2L})\Gamma(\frac{d}{2})}{\Gamma(\frac{d}{2L})\Gamma(\frac{d+1}{2})}.$ (33)
Next, we consider $\mathbb{E}[T^{\prime}(d^{\prime},L)]$. Because of the rotational symmetry of the sphere, the distribution of the inner product $Z=\langle u,v\rangle$ (for fixed $v$ and uniformly random $u$) depends only on the dimension $d^{\prime}$. It has the following density on [1,1][-1,1]:
$f_{Z}(z)=c_{d^{\prime}}(1-z^{2})^{\frac{d^{\prime}-3}{2}}$ (34)
where $c_{d^{\prime}}$ is defined as follows.
$c_{d^{\prime}}=\frac{\Gamma\left(\frac{d^{\prime}}{2}\right)}{\sqrt{\pi}\,\Gamma\left(\frac{d^{\prime}-1}{2}\right)}.$ (35)
Let $Z_{1},\dots,Z_{n}$ be i.i.d. copies of $Z=\langle u,v\rangle$. Then:
@@ Appendix C Supplementary Materials
According to Ineq. (9), there are three core terms I, II, and III, where $\text{I}=\Sigma^{L}_{i=1}{\bm{{q_{i}}^{\top}u^{i}_{e[i]}}}$, $\text{II}=A_{S}(\bm{v})$, and $\text{III}=\frac{\|\bm{w}\|^{2}/2-\tau-\bm{v}^{\top}\bm{q}}{\|\bm{e}\|}$. Here, I/II is exactly the second designed function $K_{S}^{2}$; the computation of I depends on PQ Jégou et al. (2011). III denotes the threshold for $K^{2}_{S}$ and is explained in the original paper of PEOs (see routing test Eq. (9) in Lu et al. (2024)).
@@ C.3 ANNS and Similarity Graphs
As one of the most fundamental problems, approximate nearest neighbor search (ANNS) has seen a surge of interest in recent years, leading to the development of numerous approaches across different paradigms. These include tree-based methods (Curtin et al., 2014), hashing-based methods (Andoni & Indyk, 2008; Lei et al., 2019; Andoni & Beaglehole, 2022; Andoni et al., 2015; Pham & Liu, 2022), vector quantization (VQ)-based methods (Jégou et al., 2011; Ge et al., 2014; Babenko & Lempitsky, 2016; Guo et al., 2020), learning-based methods (Gupta et al., 2022; Li et al., 2023), and graph-based methods (Malkov & Yashunin, 2020; Subramanya et al., 2019; Fu et al., 2019; 2022; Gao & Long, 2023; Peng et al., 2023; Xie et al., 2025; Yang et al., 2025).
@@
Among these, graph-based methods are widely regarded as the state-of-the-art. Currently, three main types of optimizations are used to enhance their search performance: (1) improved edge-selection strategies, (2) more effective routing techniques, and (3) quantization of raw vectors. These approaches are generally orthogonal to one another. Since the proposed structure, PEOs, belongs to the second category, we briefly introduce several highly relevant works below. TOGG-KMC (Xu et al., 2021) and HCNNG (Muñoz et al., 2019) use KD-trees to determine the direction of the query and restrict the search to points in that direction. While this estimation is computationally efficient, it results in relatively low query accuracy, limiting their improvements over HNSW (Malkov & Yashunin, 2020). FINGER (Chen et al., 2023) examines all neighbors and estimates their distances to the query. For each node, FINGER locally generates promising projection vectors to define a subspace, then uses collision counting which is similar to SimHash to approximate distances within each visited subspace. Learn-to-Route (Baranchuk et al., 2019) learns a routing function using auxiliary representations that guide optimal navigation from the starting vertex to the nearest neighbor. Recently, Lu et al. (2024) proposed PEOs, which leverages space partitioning and random projection techniques to estimate a random variable representing the angle between each neighbor and the query vector. By aggregating projection information from multiple subspaces, PEOs substantially reduces the variance of this estimated distribution, significantly improving query accuracy.
@@ C.4 Algorithms Related to KS1 and KS2
Alg. 4 and Alg. 5 are used to compare the probing accuracies of CEOs and KS1, while Alg. 6 presents the graph-based search equipped with the KS2 test.
@@
Input: $q$ is the query; $\mathcal{D}$ is the dataset; $\mathcal{I}$ is the index structure returned by Alg. 4; $k$ denotes the value of top-$k$; $s_{0}$ is the number of scanned top projection vectors; and #probe denotes the number of probed points for each projection vector

### v3

@@ 7.1 Comparison with CEOs
As demonstrated in Sec. 6.1, the results of CEOs are directly used to accelerate other similarity search processes (Pham and Liu, 2022; Xu and Pham, 2024). In this context, we focus solely on the improvement of CEOs itself. We show that KS1, equipped with the structures $S_{\text{sym}}(m,1)$ and $S_{\text{pol}}(m,1)$, can slightly outperform CEOs($m$) on the original task of CEOs, that is, $k$-MIPS, where $m$ denotes the number of projection vectors and was set to 2048, following the standard configuration of the original CEOs. Since the only difference among the compared approaches is the configuration of the projection vectors, we use a unified algorithm (see construction of projection structure Alg. 5 and MIPS query processing Alg. 6 in Appendix) with the configuration of projection vectors as an input to compare their recall rates. From the results in Table. 1, we observe that: (1) in most cases, KS1 with the two proposed structures achieves slightly better performance than CEOs, supporting our claim that a smaller reference angle yields a more accurate estimation, and (2) $S_{\text{pol}}$ generally achieves a higher recall rate than $S_{\text{sym}}$, verifying that a configuration closer to the best covering yields better performance. Since the gain of KS1 over CEOs originates from only one difference, that is, the distribution of the generated projected vectors (Gaussian vs. a more diverse distribution on the unit sphere), this result supports our claim that the Gaussian distribution is suboptimal.
@@ 7.2 ANNS Performance
We chose ScaNN (Guo et al., 2020), HNSW (Malkov and Yashunin, 2020), and HNSW+PEOs (Lu et al., 2024) as baselines, where ScaNN is a state-of-the-art quantization-based approach that performs better than IVFPQFS, and HNSW+PEOs (Lu et al., 2024) is a state-of-the-art graph-based approach that outperforms FINGER (Chen et al., 2023) and Glass (Zilliz, 2023). Like HNSW+PEOs, KS2 is implemented on HNSW, dubbed HNSW+KS2. DiskANN (Subramanya et al., 2019) is excluded because it is orthogonal to KS2 and focuses on optimization for external storage. ADSampling (Gao and Long, 2023) is excluded because it is designed for a non-SIMD environment. The parameter settings of all compared approaches and additional experimental results can be found in Appendix D.
@@
(2) Query performance. From the results in Fig. 1, we make the following observations. (i) Except for Word, HNSW+KS2 achieves the best performance among all compared methods. In particular, HNSW+KS2 accelerates HNSW by a factor of 2.5 to 3, and is 1.1 to 1.3 times faster than HNSW+PEOs, demonstrating the superiority of KS2 over PEOs. (ii) Compared with ScaNN, the advantage of HNSW+KS2 is especially evident in the recall region below 85%, highlighting the effectiveness of the routing test. On the other hand, in the high-recall region for Word, ScaNN outperforms HNSW+KS2 due to the connectivity issues of HNSW.
@@ Acknowledgements
This work is supported by JSPS Kakenhi JP22H03594, JP23K17456, JP23K25157, JP23K28096, JP25H01117, JP25K21272, and JST CREST JPMJCR22M2.
@@ Appendix A Notations
$\|\bm{e}\|$ Length of edge $\bm{e}$
$\tau$ Threshold determined by the result list of graph search
@@
$\bm{u}^{i}_{j}$ $j$-th sub-vector of the $i$-th subspace of $\bm{u}$, where $1\leq i\leq L$ and $1\leq j\leq m$
@@ Appendix B Proof of Lemmas
If $\bm{u}$ is a vector randomly drawn from $C$, the CDF of $\left\langle\bm{v},\bm{u}\right\rangle$ is $I_{t}(\frac{d-2}{2},\frac{d-2}{2})$, where $t=\frac{1}{2}+\frac{x-\cos\phi\cos\psi}{2\sin\phi\sin\psi}$.
@@
We consider the following four cases based on the value of $\left\langle\bm{q},\bm{v}\right\rangle$.
Case 1: $0<\left\langle\bm{q},\bm{v}\right\rangle<1$.
@@
Claim: Let $O,A,B,C$ be four points in $\mathbb{R}^{d}$. $\overrightarrow{OA}$ is perpendicular to $\overrightarrow{AB}$, and $\overrightarrow{OA}$ is perpendicular to $\overrightarrow{AC}$. The angle between $\overrightarrow{OA}$ and $\overrightarrow{OC}$ is $\psi$, and the angle between $\overrightarrow{OA}$ and $\overrightarrow{OB}$ is $\phi$. If the angle between $\overrightarrow{AB}$ and $\overrightarrow{AC}$ is $\alpha$, then $\cos\beta$, where $\beta$ denotes the angle between $\overrightarrow{OB}$ and $\overrightarrow{OC}$, can be expressed as follows.
@@
Case 2: $\left\langle\bm{q},\bm{v}\right\rangle=0$.
@@
Case 3: $1<\left\langle\bm{q},\bm{v}\right\rangle<0$.
@@
Case 4: $\left\langle\bm{q},\bm{v}\right\rangle=1$ or $\left\langle\bm{q},\bm{v}\right\rangle=-1$.
In this case, $\left\langle\bm{q},\bm{u}\right\rangle=\pm\cos\psi$, and the conclusion is trivial.
@@
2   Generate $m$ points randomly and independently on $\mathbb{S}^{d^{\prime}-1}$  Scale the norm of all $m$ points in this iteration to $1/\sqrt{L}$ and collect the vectors after scaling
@@ B.3 Estimation of $J(S_{\mathrm{sym}}(m,L))$
Then we only need to focus on $J(S_{\text{ran}}(m,L))$ and prove that it is equal to the RHS of Ineq. (29). Let $\bm{v}=[\bm{v_{1}},\bm{v_{2}},\cdots,\bm{v_{L}}]\in\mathbb{R}^{d}$ be a vector drawn randomly from $\mathbb{S}^{d-1}$, where $d$ is assumed to be divisible by $L$ and $d=Ld^{\prime}$. Let $r_{L}(\bm{v})$ be the subspace-normalized vector w.r.t. $\bm{v}$. That is,
@@
Let $T(d,L)=\left\langle\bm{v},r_{L}(\bm{v})\right\rangle$. For every $i$, let $\bm{u_{i}}=\bm{v_{i}}/\|\bm{v_{i}}\|\in\mathbb{S}^{d^{\prime}-1}$. Then we select $m$ vectors from $\mathbb{S}^{d^{\prime}-1}$ randomly and independently. Suppose that, among $m$ generated vectors, $\bm{w}$ is the vector having the smallest angle to $\bm{u_{i}}$, and we use $T^{\prime}(d^{\prime},m)$ to denote $\left\langle\bm{w},\bm{u_{i}}\right\rangle$. Since the choice of $\bm{w}$ for every $\bm{u_{i}}$ is independent of the value of $T(d,L)$, we have the following result:
$\mathbb{E}[A_{S}(\bm{v})]=\mathbb{E}[T(d,L)]\times\mathbb{E}[T^{\prime}(d^{\prime},m)].$ (32)
@@
$\mathbb{E}[T(d,L)]=\sqrt{L}\frac{\Gamma(\frac{d+L}{2L})\Gamma(\frac{d}{2})}{\Gamma(\frac{d}{2L})\Gamma(\frac{d+1}{2})}.$ (33)
Next, we consider $\mathbb{E}[T^{\prime}(d^{\prime},m)]$. Because of the rotational symmetry of the sphere, the distribution of the inner product $Z=\langle u,v\rangle$ (for fixed $v$ and uniformly random $u$) depends only on the dimension $d^{\prime}$. It has the following density on [1,1][-1,1]:
$f_{Z}(z)=c_{d^{\prime}}(1-z^{2})^{\frac{d^{\prime}-3}{2}}.$ (34)
Let $Z_{1},\dots,Z_{m}$ be i.i.d. copies of $Z=\langle u,v\rangle$. Then:
@@ Appendix C Supplementary Materials
According to Ineq. (9), there are three core terms I, II, and III, where $\text{I}=\sum^{L}_{i=1}{\bm{{q_{i}}^{\top}u^{i}_{e[i]}}}$, $\text{II}=A_{S}(\bm{e})$, and $\text{III}=\frac{\|\bm{w}\|^{2}/2-\tau-\bm{v}^{\top}\bm{q}}{\|\bm{e}\|}$. Here, I/II is exactly the kernel function $K_{S}^{2}$; the computation of I depends on PQ (Jégou et al., 2011). III denotes the threshold for $K^{2}_{S}$ and is explained in the original paper of PEOs (see the routing test Eq. (9) in Lu et al. (2024)).
@@ C.3 ANNS and Similarity Graphs
As one of the most fundamental problems, approximate nearest neighbor search (ANNS) has seen a surge of interest in recent years, leading to the development of numerous approaches across different paradigms. These include tree-based methods (Curtin et al., 2014), hashing-based methods (Andoni and Indyk, 2008; Lei et al., 2019; Andoni and Beaglehole, 2022; Andoni et al., 2015; Pham and Liu, 2022), vector quantization (VQ)-based methods (Jégou et al., 2011; Ge et al., 2014; Babenko and Lempitsky, 2016; Guo et al., 2020), learning-based methods (Gupta et al., 2022; Li et al., 2023), and graph-based methods (Malkov and Yashunin, 2020; Subramanya et al., 2019; Fu et al., 2019; 2022; Gao and Long, 2023; Peng et al., 2023; Xie et al., 2025; Yang et al., 2025).
@@
Among these, graph-based methods are widely regarded as the state-of-the-art. Currently, three main types of optimizations are used to enhance their search performance: (1) improved edge-selection strategies, (2) more effective routing techniques, and (3) quantization of raw vectors. These approaches are generally orthogonal to one another. Since PEOs belongs to the second category, we briefly introduce several highly relevant works below. TOGG-KMC (Xu et al., 2021) and HCNNG (Muñoz et al., 2019) use KD-trees to determine the direction of the query and restrict the search to points in that direction. While this estimation is computationally efficient, it results in relatively low query accuracy, limiting their improvements over HNSW (Malkov and Yashunin, 2020). FINGER (Chen et al., 2023) examines all neighbors and estimates their distances to the query. For each node, FINGER locally generates promising projection vectors to define a subspace, then uses collision counting which is similar to SimHash to approximate distances within each visited subspace. Learn-to-Route (Baranchuk et al., 2019) learns a routing function using auxiliary representations that guide optimal navigation from the starting vertex to the nearest neighbor. Recently, Lu et al. (2024) proposed PEOs, which leverages space partitioning and random projection techniques to estimate a random variable representing the angle between each neighbor and the query vector. By aggregating projection information from multiple subspaces, PEOs substantially reduces the variance of this estimated distribution, significantly improving query accuracy.
Recently, a related quantization method RaBitQ (Gao and Long, 2024) was proposed and used for the acceleration of similarity graphs (Gou et al., 2025). Notably, RaBitQ can be viewed as a special case of KS2: when $m=2$ and $L=d$, RaBitQ has exactly the same projection structure as KS2. On the other hand, Lu et al. (2024) shows that when $L$ is very large, an additional error will increase significantly. Thus, if the tradeoff between index size and search performance is taken into account, choosing $L=d$ may not be ideal. Meanwhile, when RaBitQ is applied to similarity graphs, the additional overhead to accommodate the index is very large, consuming about 2–4 times the size of the dataset.
Algorithm 4 Graph-based ANNS with routing
Input: Query $\bm{q}$, # results $K$, graph index $G$
Output: $K$-NN of $\bm{q}$
$R\leftarrow\emptyset$ // an ordered list of results, $|R|\leq efs$
$P\leftarrow\{\,\text{entry node }v_{0}\in G\,\}$ // a priority queue
while $P\neq\emptyset$ do
4 $v\leftarrow P.pop()$; foreach unvisited neighbor $u$ of $v$ do
5    if $|R|<efs$ then $\delta\leftarrow\infty$;
6    else $p\leftarrow R[efs]$, $\delta\leftarrow dist(\bm{p},\bm{q})$;
7    if RoutingTest($\bm{u},\bm{v},\bm{q},\delta$) = true then
8       if $dist(\bm{u},\bm{q})<\delta$ then
9          $R.push(\bm{u})$; $P.push(\bm{u})$;
10
11
12
return ($\{\,R[1],\ldots,R[K]\,\}$)
@@ C.4 Algorithms Related to KS1 and KS2
Alg. 5 and Alg. 6 are used to compare the probing accuracies of CEOs and KS1, while Alg. 7 presents the graph-based search equipped with the KS2 test. Because Alg. 6 resembles the probabilistic routing mechanism in Lu et al. (2024), we provide further clarification through Alg. 7.
In Alg. 7, the list $R$ serves as an ordered container of at most efsefs elements ($efs\geq K$). Using a graph index $G=(V,E)$ constructed over the dataset, ANNS results are obtained by traversing this graph, as summarized in Alg. 7. The search starts from an entry node $v_{0}\in G$ and maintains $R$, which holds the current best candidates, together with a priority queue $P$ that stores nodes yet to be explored. A neighbor of the current node is inserted into the priority queue whenever it is closer to $\bm{q}$ than the farthest element in $R$, or whenever $R$ has not reached its capacity. Nodes are repeatedly popped from the priority queue to explore their neighbors, and the procedure terminates once the queue becomes empty. In practice, additional optimizations—such as early stopping or pruning strategies—are commonly adopted to further accelerate the process (Malkov and Yashunin, 2020).
A straightforward implementation would compute the exact distance between $\bm{q}$ and every encountered neighbor. However, a routing test can be used to decide whether an exact distance evaluation is needed. The probabilistic routing technique (Lu et al., 2024) is one such example. Meanwhile, the proposed function $K^{2}_{S}$ naturally leads to an alternative routing criterion, given by Ineq. (9), which can also be used to determine whether a neighbor should undergo exact distance calculation.
@@
Input: $q$ is the query; $\mathcal{D}$ is the dataset; $\mathcal{I}$ is the index structure returned by Alg. 5; $k$ denotes the value of top-$k$; $s_{0}$ is the number of scanned top projection vectors; and #probe denotes the number of probed points for each projection vector